OpenStudy (anonymous):

Construct a polynomial that has zeros at -2;-1; 3; third degree; P(2) = 24.

OpenStudy (gorv):

let the equation is p(x)=ax^3+bx^2+cx+d each root will satisfy this equation p(-2)=p(-1)=p(3)=0 put x= -2 , -1 , 3 and p(2)= 24 given and solve

OpenStudy (solomonzelman):

well, when you know that number `v` is zero, the root would be written as `(x-v)` So in your case, f(x)=(x+2)(x+1)(x-3) now, you just need to expand to receive the desired function.